Observer ะž notes that two events are separated in space and time by 600 ๐‘š and 8 ร— 10โˆ’7


๐‘ . How fast


must an observer O' be moving relative to ะž in order that the events be simultaneous to O'?

1
Expert's answer
2022-02-15T10:55:47-0500

d=600m

"t=8\times10^{-7}sec"

"O" co-ordination System

"v=\frac{d}{t}"

"v=\frac{600}{8\times10^{-7}}=75\times10^{7}m/sec"

Now "O'" co -ordination system

Similarly

d'=600m

"t'=8\times10^{-7}sec"

"v'=\frac{d'}{t'}"

"v'=\frac{600}{8\times10^{-7}}=75\times10^{7}m/sec"

"v=v'"